how much pure acid should be mixed with 7 gallons of a 40% acidic solution in order to get a 80% acidic solution
Answer by mananth(16946)

--- percent ---------------- Amount
Acid I 40 ---------------- 7 gallons
Pure Acid 100 ---------------- x gallons
Mixture 80 ---------------- 7 + x gallons

40 * 7 + 100 x = 80 ( 7 + x )
280 + 100 x = 560 + 80 x
100 x -80 x = -280 + 560
20 x = 280
/ 20
x = 14 gallons Pure Acid

m.ananth@hotmail.ca